jakfishman:

Fill in the table below to see if the Pythagorean Theorem is true for other right triangles. First use the numbers given in the first two rows. Then try two right triangles of your own. (Note: The value of c in the Gizmo is rounded, so you may find small rounding errors.) a b c a2 b2 a2 + b2 c2 3 10 7 7

jakfishman:

aidenj:

well what do u think so far

jakfishman:

i have to add a and b to make c?

aidenj:

do u know how to do the pythagorean therom

jakfishman:

not rrly

jakfishman:

she trys ive talked in dms with her shes doing school

surjithayer:



surjithayer:

in the diagram ABC is a right angled triangle with right angle at C. then according to pythagoras theorem sum of squares of two sides=square of hypotenuse. or here \[a^2+b^2=c^2\]

surjithayer:

\[a^2=3^2=9\] \[b^2=10^2=109\] \[a^2+b^2=9+100=109\] \[c^2=109\] \[c=\sqrt{109}\] similarly solve other problems.